The pollutant sulphur dioxide can removed from the air by reaction with calcium carbonate in the presence of oxygen. What mass of calcium carbonate is needed to remove 1 tonne of sulphur dioxide? 

2CaCO3 + 2SO2 + O2 → 2CaSO4 + 2CO2

1
Expert's answer
2022-01-18T06:08:02-0500

Molar mass of SO2 = 64 .066g


Molar mas of CaCO3 = 100.0869


Mole ratio SO2:CaCO3=1:1


1 tonne = "1\u00d710^6g"


Let m= tonnes of CaCO3


"\frac{1×10^6}{64.066}=\frac{m×10^6}{100.0869}"


"m= 1.562\>\> tonnes"